# Painless Activation Steering: An Automated, Lightweight Approach for Post-Training Large Language Models

### v2

@@ 3.2 Causal Steering Effect on behavior & Intelligence-Oriented Tasks
We find that all three proposed steering methods yield statistically significant improvements on every behavior task. The causal steering effects of all PAS methods are positive and meaningful. iPASwo is the most effective method for Bias ($\Delta=0.101$), PASf is most effective for Morality ($\Delta=0.089$), and PASf and iPASa are for Alignment ($\Delta=0.425$), with $p<0.005$ throughout. Moreover, we find that the two introspective strategies outperform PASf for 13 out of the 15 tasks, highlighting the effectiveness of introspective strategies, in which the steering vectors are constructed solely from the model’s own mistakes. LABEL:{tab:unsteered_acc_all} reports the accuracy of the unsteered raw model and Table 6 reports full point estimates, confidence intervals, and $p$-values of the causal steering effect. [^4] Throughout the paper, $\textit{p}=0.00$ indicates that the $p$-value is smaller than $5\times 10^{-3}$. Taken together, each PAS strategy significantly improves performance across a broad set of behavior tasks.
Result: PAS Has Mixed Effectiveness on Intelligence-Oriented Tasks
We find that the effectiveness of PAS on intelligence-oriented tasks is mixed. For Deepseek-R1-Distill-7B, PAS yields statistically significant improvements on all three evaluation dimensions, with PASf achieving the strongest gains (e.g., $\Delta=0.072$ on OpenBookQA, $\Delta=0.066$ on ARC Challenge, $\Delta=0.026$ on LSAT, all with $p<0.001$). In contrast, for Nous-Hermes-2-Mistral and Llama-3.1-8B-Instruct, only negligible or inconsistent improvements are achieved and most are statistically insignificant. Taken together, PAS does not consistently perform well on intelligence-oriented tasks, suggesting that PAS cannot yet substitute weight-based post-training methods on tasks requiring specialized knowledge or advanced reasoning capacity. Table 3 reports full point estimates, confidence intervals, and $p$-values of the causal steering effects on intelligence-oriented tasks. We further analyze this limitation and discuss potential improvements in Section 4.
3.3 Causal Steering Effects at Various Steer Targets
@@
There are several possible steer targets to where the hook may be attached: the residual stream between sub-modules (residual); the multi-head self-attention module (self_attn); the LayerNorm applied after attention and before MLP (post_attn); and the feedforward block, capturing the nonlinear transformation (mlp). Most experiments in this paper use residual as the default steer_targ, consistent with standard practice in the literature. We experiment with the three alternative steer targets to investigate whether PAS is architecturally sensitive. [^5] It is possible to construct the steering vector $a^{*}$ from a steer_targ of layer $\ell$ and inject it to a different $(\text{{steer\_targ}}^{\prime},\ell^{\prime})$ pair, but we abandoned this line of inquiry after unpromising preliminary results.
Result: Alternative Steer Targets also Work, Albeit Less Effectively Than Residual
We find that PAS is architecturally robust, consistently generates statistically significant causal steering effects across steer targets. Averaging over the three steering strategies, we observe consistent improvements across all evaluation dimensions at each steering location. For self_attn (Table 7), the average accuracy improvement is $\Delta=0.051$ for Bias, $\Delta=0.044$ for Morality, and $\Delta=0.328$ for Alignment. For post_attn (Table 8), the improvements are $\Delta=0.021$ for Bias, $\Delta=0.031$ for Morality, and $\Delta=0.301$ for Alignment. For mlp (Table 9), the gains are $\Delta=0.038$ for Bias, $\Delta=0.039$ for Morality, and $\Delta=0.313$ for Alignment.
3.4 Causal Steering Effects on Top of ICL
@@
How does PAS compare with ICL? We extend our experiment to the ICL setting. Starting from $M$, we provide 10 in-context exemplars drawn from the model’s incorrect answers on the training split. We re-evaluate the model on the training set under ICL and construct the steering vectors based on its performance. Finally, we assess the ICL model with and without steering. In other words, we directly compare “ICL-only” with “ICL+PAS,” testing whether steering vectors constructed from errors after ICL can further improve performance on the held-out test set.
Result: PAS Provides Additional Gains on Top of ICL
@@
We find that applying PAS on top of ICL yields additional gains beyond ICL alone (Table 10). iPASa, iPASwo, and PASf improve performance on Bias by 3.0%, 3.2%, and 2.4%, on Morality by 3.9%, 4.3%, and 2.0%, and on Alignment by 16.1%, 16.7%, and 18.1%.
3.5 Causal Steering Effects on Top of SFT
Result: PAS Outperforms SFT in Causal Steering Effects on TruthfulQA
We find that, on TruthfulQA, running PAS on top of an SFT-trained model ($\overline{M^{\prime}}$) beats SFT alone ($M^{\prime}$). More surprising, the performance of the PAS-trained base model ($\overline{M}$) is statistically indistinguishable from that of a model trained with both PAS and SFT ($\overline{M^{\prime}}$), implying that once PAS is applied, SFT provides no additional benefit.
To support the former statement, we conduct Hypothesis Test A, where PAS applied on top of SFT yields an average improvement of $0.27$ for PASf and 0.140.140.150.15 for the introspective variants, all with $p<0.01$, while the benefits of SFT over the base model are only $0.09$ (Hypothesis Test B). Hypothesis Test C confirms that the PAS-trained model $\overline{M}$ wins against the SFT-trained model $M^{\prime}$ decisively: $0.22$ for PASf and 0.140.140.150.15 for the introspective methods ($p<0.01$).
@@ 3.6 Catastrophic Forgetting
A post-training method is useful if, in addition to improving the target evaluation dimension (i.e., $\mathbb{E}\left[Y_{k}(\overline{M})-Y_{k}(M)\right]>\epsilon_{k}$), it also maintains performance on the control evaluation dimensions (i.e., $\mathbb{E}\left[Y_{\phi}(\overline{M})-Y_{\phi}(M)\right]>-\epsilon_{\phi}$ for $\phi\in\Phi$). If performance on the control evaluation dimension decays significantly, we say that the model suffers from catastrophic forgetting. To quantify catastrophic forgetting, we evaluate $\overline{M}$ on out-of-domain tasks from MMLU (Hendrycks et al., 2020b) after applying PAS. For each source task $k$ and method $m$, we define $\Delta\mathrm{MMLU}(k,m)$ as the average change in MMLU accuracy across the 57 subjects, relative to the unsteered model.
Result: PAS Generally Does Not Cause Catastrophic Forgetting
@@
For most tasks, the point estimates of the catastrophic forgetting effects are negligible—that is, steering does not significantly degrade control evaluation dimensions (see Table 5). The two exceptions are Sycophancy and TruthfulQA, which show substantial drops in accuracy. Averaged over the 3 strategies and models, Sycophancy drops by $\Delta=-0.21$ and TruthfulQA drops by $\Delta=-0.13$. However, further analysis revealed that these large drops were caused by high steering strengths (up to 32 for Sycophancy and 8 for TruthfulQA). When we restricted the strength range to 0–5, the average catastrophic effect across the three models for both tasks decreased to 9%. This finding supports our practical recommendation of setting the steering strength to 1 (see Section 3.7). In practice, since PAS can be easily enabled or disabled, we recommend turning it off when the LM is not used for target tasks.
3.7 Hyperparameter and Sample Size Sensitivity Analysis
@@
We conduct a grid search over steering layers and steering strengths using the validation split to find the best hyperparameters for PAS. We also run PAS with a range of sample sizes (from small data (n-train= 12) to large data (n-train=2400) to investigate the influence of sample size.
Result: Moderate Steering Strength Applied at Middle Layer Works the Best
@@
Fig. 3 reports how iPASwo’s accuracy varies with these choices on two representative tasks across three models. (More details are reported in Appendix A.) Across nearly all 15 behavior tasks and three models (see Figs. 4 and 5) we observe a consistent pattern. First, iPASwo performs best when the steering vector is injected in the middle layers of the transformer (around layer 14 in 32-layer LMs), with gains declining in shallower or deeper layers. Second, performance varies with steering strength in a concave manner: very small coefficients yield limited gains, excessively large ones degrade accuracy, and a moderate strength ($\lambda\approx 1$) gives the strongest improvements. Remarkably, accuracy is independent of the steering vector’s norm and depends only on steering strength, suggesting that the scaling relative to the representation space—rather than absolute vector magnitude—governs the effectiveness of activation steering.
Result: PAS Is Effective In Small Data Regime
@@
Across tasks, accuracy generally increased with more training data, but only moderately. This shows that PAS is relatively insensitive to sample size and remains effective even in small data regime. More details are reported in Appendix D.
Default Hyperparameter Recommendations
Based on our ablation experiments, we recommend the following default PAS hyperparameters: target the middle third of the model’s layers (for example, layer 14 or 16 in a 32-layer LM), use the residual stream as the steer target, set the steering strength around $\lambda\approx 1$, and use more data if available (though 2000 typically suffices).
4 Follow-Up Work and Limitations
@@
Theoretical Analysis of Limitations & Multi-Layer Activation Steering
PAS is ineffective on intelligence-oriented tasks, and we hope to understand why. Here is a possible angle of analysis. Suppose injection occurs in the residual stream at layer $\ell$. Let the unsteered activation be $z^{(\ell)}=W^{(\ell)}h^{(\ell-1)}+b^{(\ell)}$; after steering, it becomes
$\overline{z}^{(\ell)}=W^{(\ell)}(h^{(\ell-1)}+\lambda a^{*})+b^{(\ell)}=W^{(\ell)}h^{(\ell-1)}+\big(b^{(\ell)}+\lambda W^{(\ell)}a^{*}\big).$
@@
This invites follow-up work to characterize the Pareto frontier between the strong causal steering effects of full-parameter RL and the speed and convenience of PAS. It also motivates the extension to Multi-Layer Activation Steering (applying AS across layers or combining several task vectors).
Alternative Extraction Methods
Beyond computing the mean activation difference without post-processing or dimensionality reduction, as we do here for simplicity, we can explore alternative vector extraction methods, such as whitening (Kalapos & Gyires-Tóth, 2024), probing (O’Neill et al., 2025; Goldowsky-Dill et al., 2025), or separating with linear discriminants (Fisher, 1936).
Alternative Applications
Beyond behavior tasks, we are exploring the use of AS for diverse purposes, such as Compressive Activation Steering (where vectors contrast is used to encode and encrypt private information), Stylistic Activation Steering (where vectors are designed to induce writing styles of specific authors), Forgetting Activation Steering (where vectors suppress targeted knowledge such as all events since COVID-19), and Speculative Steering (where AS is turned on selectively based on a classifier layer (Lee et al., 2025) or a teacher model (Leviathan et al., 2023)).
5 Conclusion
We introduce Painless Activation Steering (PAS), a fully automated approach that makes activation steering fast, human-independent, and practical. Across three open-weight models and 15 diverse evaluation dimensions, PAS (especially its introspective variants) delivers consistent gains on behavior tasks. We provide statistically significant evidence that, over a wide range of settings, PAS can be a cheaper and lighter alternative to weight- and prompt-based post-training, offering new opportunities for modular and adaptive control of LMs. By systematically characterizing where PAS helps, hurts, and complements existing approaches, we hope to establish activation steering as a practical, human-independent, and automation-friendly recipe for post-training, well-suited for non-intelligence-oriented personalization and customization. We view PAS as a promising foundation for future work on fast and flexible LM post-training methods. We invite researchers and practitioners to explore the potential of PAS in their behavioral post-training applications.
Reproducibility Statement
@@ 5 Conclusion
We provide a GitHub repository at https://github.com/Sasha-Cui/Painless_Activation_Steering that contains all information necessary to reproduce the experimental results. All open-source LMs and datasets used are publicly available on Hugging Face.
Ethics Statement
@@
LM Usage Statement
    Please check the grammar of my draft.  Thanks.
    Let’s look at the neural activation editing based steering
    methods for LLMs such as contractive activation addition
    and the transluce monitor.
    1.  A comparative analysis.  I want to know what the current
    limitations of understandings are regarding these neural
    steering methods.
    2.ΨI want to know if I can make a model nicer and more
    logical and more interested in talking about the bible at
    the same time.
    3.ΨThe focus is on LLMs.
      Please critique this paper in an honest, direct, and detailed
      fashion. Be accurate and think deeply. Do online searches
      to find related work. Please check the math carefully. Take
      your time.
Acknowledgments
We thank Chen Guang, Anurag Kahyap, Jasjeet Sekhon, and Samuel Soo for helpful discussions. We thank Jasjeet Sekhon, Yale Center for Research Computing (YCRC), and Massachusetts High Performance Computing Center (MGHPCC) for funding and providing computational resources.
Appendix A Hyperparameter Analysis
We present the plots showing how validation split accuracy varies with the hyperparameter across 15 behavior tasks and 3 LMs. We vary the layer from $8,9,\ldots,25$ and the steering strength from $0.25,0.5,0.75,1.0,4.0,7.0,10.0,\ldots,32.0$. Fig. 4 and Fig. 5 report how iPASwo’s validation accuracy varies with steering strength and layers across 15 behavior tasks. The results are obtained by taking the maximum over another hyperparameter.

### v3

@@ 3.2 Steering Effect on Behavior & Intelligence Tasks
Result 1: PAS reliably improves behavior tasks. All three PAS variants yield statistically significant improvements on every behavior task and generally outperform the Contrastive Activation Addition (CAA) baseline (Panickssery et al., 2024) (Figure 2). The strongest gains are domain-dependent: iPASwo performs best on Bias, PASf performs best on Morality, and PASf/iPASa perform best on Alignment. Moreover, the two introspective strategies outperform PASf on 13 of the 15 behavior tasks, suggesting that vectors constructed from a model’s own mistakes are especially effective. Full unsteered accuracies and task-level estimates are reported in Tables 2 and 5.444Throughout the paper, $\textit{p}=0.00$ indicates that the $p$-value is smaller than $5\times 10^{-3}$.
Result 2: PAS does not consistently improve intelligence-oriented tasks. PAS has mixed effects on OpenBookQA, ARC Challenge, and LSAT. It improves all three intelligence-oriented tasks for DeepSeek-R1-Distill-7B, but yields negligible or inconsistent gains for Nous-Hermes-2-Mistral and Llama-3.1-8B-Instruct. Thus, PAS is better understood as a lightweight behavioral post-training method than as a substitute for weight-based training on knowledge- or reasoning-intensive tasks. Full estimates are reported in Table 3.
3.3 Steering Effects at Various Steer Targets
@@
There are several possible steer targets to where the hook may be attached: the residual stream between sub-modules (residual); the multi-head self-attention module (self_attn); the LayerNorm applied after attention and before MLP (post_attn); and the feedforward block, capturing the nonlinear transformation (mlp). Most experiments in this paper use residual as the default st, consistent with standard practice in the literature. We experiment with the three alternative steer targets to investigate whether PAS is architecturally sensitive. [^5] It is possible to construct the steering vector $a^{*}$ from a st of layer $\ell$ and inject it to a different $(\text{{st}}^{\prime},\ell^{\prime})$ pair, but we abandoned this line of inquiry after unpromising preliminary results.
Result 3: PAS works across steering targets, but is strongest in the residual stream. PAS remains effective when applied to self_attn, post_attn, or mlp, indicating that the method is not tied to a single architectural location. However, the residual stream remains the strongest and most stable default target. Full estimates for alternative targets are reported in Tables 6, 7 and 8.
3.4 Additional Steering Effects on Top of ICL
@@
How does PAS compare with ICL? We extend our experiment to the ICL setting. Starting from $M$, we provide 10 in-context exemplars drawn from the model’s incorrect answers on the training split. We re-evaluate the model on the training set under ICL and construct the steering vectors based on its performance. Finally, we assess the ICL model with and without steering. In short, we compare “ICL-only,” “PAS-only,” and “ICL+PAS.” We particularly test if steering vectors constructed from errors after ICL can further improve performance on the held-out test set.
@@
Result 4: PAS complements ICL. We find that PAS alone is not consistently better than ICL alone (Table 10). However, applying PAS on top of ICL yields additional gains beyond ICL alone (Table 9). iPASa, iPASwo, and PASf improve performance on Bias by 3.0%, 3.2%, and 2.4%, on Morality by 3.9%, 4.3%, and 2.0%, and on Alignment by 16.1%, 16.7%, and 18.1%.
@@ 3.5 Catastrophic Forgetting
A post-training method is useful if, in addition to improving the target evaluation dimension,
$\mathbb{E}\left[Y_{k}(\overline{M})-Y_{k}(M)\right]>\epsilon_{k},$
it also maintains performance on the control evaluation dimensions,
$\mathbb{E}\left[Y_{\phi}(\overline{M})-Y_{\phi}(M)\right]>-\epsilon_{\phi},\quad\phi\in\Phi.$
@@
If performance on the control evaluation dimension decays significantly, we say that the model suffers from catastrophic forgetting. To quantify catastrophic forgetting, we evaluate $\overline{M}$ on out-of-domain tasks from MMLU (Hendrycks et al., 2020b) after applying PAS. For each source task $k$ and method $m$, we define $\Delta\mathrm{MMLU}(k,m)$ as the average change in MMLU accuracy across the 57 subjects, relative to the unsteered model.
@@
Result 5: PAS usually avoids catastrophic forgetting under moderate steering strength. For most tasks, the point estimates of the catastrophic forgetting effects are negligible—that is, steering does not significantly degrade control evaluation dimensions (see Table 4). The two exceptions are Sycophancy and TruthfulQA, which show substantial drops in accuracy. Averaged over the 3 strategies and models, Sycophancy drops by $\Delta=-0.21$ and TruthfulQA drops by $\Delta=-0.13$. However, further analysis revealed that these large drops were caused by high steering strengths (up to 32 for Sycophancy and 8 for TruthfulQA). When we restricted the strength range to 0–5, the average catastrophic effect across the three models for both tasks decreased to 9%. This finding supports our practical recommendation of setting the steering strength to 1 (see Section 3.6). In practice, since PAS can be easily enabled or disabled, we recommend turning it off when the LM is not used for target tasks.
3.6 Hyperparameter and Sample Size Sensitivity
@@
We conduct a grid search over steering layers and steering strengths using the validation split to find the best hyperparameters for PAS. We also run PAS with a range of sample sizes, from n-train=12 to n-train=2400, to study the effect of sample size.
@@
Result 6: PAS works best with middle-layer. Figure 3 reports how iPASwo’s accuracy varies with these choices on two representative tasks across three models. (More details are reported in Appendix C.) Across nearly all 15 behavior tasks and three models (see Figures 4 and 5) we observe a consistent pattern. First, iPASwo performs best when the steering vector is injected in the middle layers of the transformer (around layer 14 in 32-layer LMs), with gains declining in shallower or deeper layers. Second, performance varies with steering strength in a concave manner: very small coefficients yield limited gains, excessively large ones degrade accuracy, and a moderate strength ($\lambda\approx 1$) gives the strongest improvements. Remarkably, accuracy is independent of the steering vector’s norm and depends only on steering strength, suggesting that the scaling relative to representation space, not absolute vector magnitude, governs the effectiveness of AS.
@@
PAS works best with moderate-strength steering. Across tasks, accuracy generally increased with more training data, but only moderately. This shows that PAS is relatively insensitive to sample size and remains effective even in small data regime. More details are reported in Appendix G.
Default Hyperparameter Recommendations Based on our ablation experiments, we recommend the following default PAS hyperparameters: target the middle third of the model’s layers (for example, layer 14 or 16 in a 32-layer LM), use the residual stream as the steer target, set the steering strength around $\lambda\approx 1$, and use more data if available (though 2000 typically suffices). Intervening on the residual stream is standard in the activation-steering literature, as it aggregates contributions from attention and MLP and directly feeds into the next layer. Our bias-shift view explicitly treats PAS as modifying the bias in the residual pathway at layer $\ell$, which is the most direct, architecture-agnostic location for a first-class steering primitive. Injecting in the middle third of layers is consistent with multiple independent observations that high-level semantic features (bias, sycophancy, moral framing) emerge in mid-late layers rather than at the input or output extremes. One possible theoretical explanation is that the latter layers capture the letter choices whereas the initial layers are still processing the problem statement. The conceptual reasoning happens in the middle layers. Our recommendation on using moderate strength is exactly what one would expect if PAS is traversing a “semantically meaningful direction” in representation space: too little movement has little effect, too much overrides useful context.
3.7 Steering Effects on Top of SFT
We examine how PAS compares with SFT. We choose Vicuna-7B as the base model $M$ and a variant that had been SFT-trained on TruthfulQA as $M^{\prime}$ (Yang et al., 2024). We conduct various hypothesis tests and find that running PAS on top of an SFT-trained model ($\overline{M^{\prime}}$) beats SFT alone ($M^{\prime}$). More surprising, the performance of the PAS-trained base model ($\overline{M}$) is statistically indistinguishable from that of a model trained with both PAS and SFT ($\overline{M^{\prime}}$), implying that once PAS is applied, SFT provides no additional benefit. Appendix I contains full details.
3.8 Open-Ended Generation
We also evaluate PAS on free-form generation by removing the multiple-choice options from the 10 bias tasks and grading answers with an external GPT-4o judge. Although accuracy drops relative to the MCQ setting (expectedly so, since the steering vectors were not adaptively built for this setting), PAS still provides consistent gains across all benchmarks, outperforming the unsteered models. Appendix H contains full details.
4 Conclusion
We introduce Painless Activation Steering (PAS), a fully automated approach that makes activation steering fast, human-independent, and practical. Across three open-weight models and 15 diverse evaluation dimensions, introspective PAS delivers consistent gains on behavior tasks. We provide statistically significant evidence that, over a wide range of settings, iPAS can be a cheaper and lighter alternative to weight- and prompt-based post-training, offering new opportunities for modular and adaptive control of LMs. By systematically characterizing where PAS helps, hurts, and complements existing approaches, we hope to establish activation steering as a practical, human-independent, and automation-friendly recipe for post-training, well-suited for non-intelligence-oriented personalization and customization. We view PAS as a promising foundation for future work on fast and flexible LM post-training methods. We invite researchers and practitioners to explore the potential of PAS in their behavioral post-training applications.
5 Limitations
@@
Theoretical Analysis of Limitations & Multi-Layer Activation Steering PAS is ineffective on intelligence-oriented tasks, and we hope to understand why. Here is a possible angle of analysis. Suppose injection occurs in the residual stream at layer $\ell$. Let the unsteered activation be $z^{(\ell)}=W^{(\ell)}h^{(\ell-1)}+b^{(\ell)}$; after steering, it becomes
$\overline{z}^{(\ell)}=W^{(\ell)}h^{(\ell-1)}+\big(b^{(\ell)}+\lambda W^{(\ell)}a^{*}\big).$
@@
This invites follow-up work to characterize the Pareto frontier between the strong steering effects of full-parameter RL and the speed and convenience of PAS. It also motivates the extension to Multi-Layer Activation Steering (applying AS across layers or combining several task vectors).
Alternative Extraction Methods Beyond computing the mean activation difference without post-processing or dimensionality reduction, as we do here for simplicity, we can explore alternative vector extraction methods, such as whitening (Kalapos and Gyires-Tóth, 2024), probing (O’Neill et al., 2025; Goldowsky-Dill et al., 2025), or separating with linear discriminants (Fisher, 1936).
Mixed Effectiveness on Intelligence and Knowledge Tasks Our systematic evaluation shows that PAS provides no reliable improvements on intelligence- and knowledge-oriented tasks such as ARC and LSAT. We view making these failures explicit as a contribution: prior AS work has not, to our knowledge, publicly documented the substantial limitations of AS on reasoning and factual evaluation settings. We hope to surface these limitations directly and steer future research away from unproductive directions.
6 Acknowledgments
We thank Chen Guang, Anurag Kahyap, Jasjeet Sekhon, and Samuel Soo for helpful discussions. We thank Jasjeet Sekhon, Yale Center for Research Computing (YCRC), Massachusetts High Performance Computing Center (MGHPCC), and Voltage Park for funding and providing computational resources, and Thomas Langford, Kaylea Nelson, Kathleen McKiernan, Aya Nawano, Michael Rothberg, John Montelus for technical assistance.
Appendix A Reproducibility Statement
@@
We provide a GitHub repository at https://github.com/Sasha-Cui/Painless_Activation_Steering/ that contains all information necessary to reproduce the experimental results. All open-source LMs and datasets used are publicly available on Hugging Face.
Appendix B Potential Risks
@@
Appendix C Hyperparameter Analysis
We present the plots showing how validation split accuracy varies with the hyperparameter across 15 behavior tasks and 3 LMs. We vary the layer from $8,9,\ldots,25$ and the steering strength from $0.25,0.5,0.75,1.0,4.0,7.0,10.0,\ldots,32.0$. Figure 4 and Figure 5 report how iPASwo’s validation accuracy varies with steering strength and layers across 15 behavior tasks. The results are obtained by taking the maximum over another hyperparameter.